# Time Weaver: A Conditional Time Series Generation Model

### v1

@@ 5 Joint Frechet Time Series Distance
In J-FTSD, we compute the FD between the real and generated joint distributions of time series and the paired metadata. Consider samples from a real data distribution indicated by $D_{r}=\left\{(x_{1}^{r},c_{1}),\ldots,(x_{n}^{r},c_{n})\right\}$, where $x_{i}^{r}$ indicates the time series and $c_{i}$ indicates the paired metadata. We denote the dataset of generated time series and the corresponding metadata as $D_{g}=\left\{(x_{1}^{g},c_{1}),\ldots,(x_{n}^{g},c_{n})\right\}$, where $x_{i}^{g}=G(c_{i})\ \forall i\in[1,n]$, and $G$ denotes any arbitrary conditional generation model defined in Sec. 3. First, similar to the FID and FJD computations, we project the time series and the paired metadata into a lower-dimensional embedding space using $\phi_{\textnormal{time}}(\cdot):\mathbb{R}^{L\times F}\rightarrow\mathbb{R}^{d_{\textnormal{emb}}}$ and $\phi_{\textnormal{meta}}(\cdot):\mathbb{R}^{L\times K}\rightarrow\mathbb{R}^{d_{\textnormal{emb}}}$ as respective feature extractors, where $d_{\textnormal{emb}}$ is the size of the embedding. We concatenate these time series and metadata embeddings to create a joint embedding space. We then calculate the FD over the joint embedding space. As such J-FTSD is formally defined as:
@@
$z^{d}_{i}$ $=\phi_{\textnormal{time}}(x^{d}_{i})\oplus\phi_{\textnormal{meta}}(c_{i})\quad\quad\forall i:(x_{i}^{d},c_{i})\in D_{d},$
@@
In essence, J-FTSD computes the FD between the Gaussian approximations of the real and generated joint embedding datasets. In Eq. 6, $\mu_{z^{r}}$, $\mu_{z^{g}}$ and $\Sigma_{z^{r}}$, $\Sigma_{z^{g}}$ are the mean and the variance of the Gaussian approximation of the real and generated joint embedding dataset respectively.
@@ 12 Stations, 5 years, 12 months,
Training Feature Extractors: Now, we describe our approach to obtain the feature extractors $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$. As explained in Sec. 2, DeVries et al. (2019) suggest using separate encoders for data samples and conditions. However, they only deal with a specific type of condition, and this naturally poses a problem for a straightforward extension of their approach to our case, where the metadata could be any arbitrary combination of categorical, continuous, and time-varying features. As such, we propose a novel approach to train the feature extractors $\phi_{\textnormal{meta}}$ and $\phi_{\textnormal{time}}$ specific to the time series domain. We jointly train $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$ with contrastive learning to better capture the joint distribution of the time series and paired metadata, as contrastive learning is a commonly used method to map data coming from various modalities into a shared latent space (Yuan et al., 2021; Zhang et al., 2022; Ramesh et al., 2022).
Figure 4: J-FTSD metric correctly penalizes the conditional time series data distribution. A good metric should penalize the conditional generation approaches for not being specific to the metadata and deviating from real time series data distribution. As such, we compare the sensitivity of different distance metrics under various synthetic disturbances on the Air Quality dataset (starting from the left); we add Gaussian noise, warp, impute, and randomly change the metadata of the time series samples. We clearly show that as the amount of perturbation increases, our J-FTSD metric (in red) shows the highest sensitivity, correctly capturing the dissimilarities between the perturbed and the original datasets, while the other metrics remain unchanged or show lower sensitivity.
@@
0:  Time series feature extractor $\phi_{\textnormal{time}}$, Metadata feature extractor $\phi_{\textnormal{meta}}$, Time series batch $X_{\textnormal{batch}}$, Paired Metadata batch $C_{\textnormal{batch}}$, Number of patches $N_{\textnormal{patch}}$, Patch length $L_{\textnormal{patch}}$, Batch size $N_{\textnormal{batch}}$.
@@
1:  Randomly select $N_{\textnormal{patch}}$ patches of length $L_{\textnormal{patch}}$ from each sample in $X_{\textnormal{batch}}$ and $C_{\textnormal{batch}}$ to generate $X_{\textnormal{batch}}^{\textnormal{patch}}$ and $C_{\textnormal{batch}}^{\textnormal{patch}}$.
@@
2:  Obtain the time series and metadata embedding - $\phi_{\textnormal{time}}(X_{\textnormal{batch}}^{\textnormal{patch}})$ and $\phi_{\textnormal{meta}}(C_{\textnormal{batch}}^{\textnormal{patch}})$ respectively.
3:  Obtain the logits - $\phi_{\textnormal{time}}(X_{\textnormal{batch}}^{\textnormal{patch}})^{T}\phi_{\textnormal{meta}}(C_{\textnormal{batch}}^{\textnormal{patch}})$.
4:  Define the labels - $[0,1,2,\dots,N_{\textnormal{batch}}\times N_{\textnormal{patch}}-1]$.
5:  Compute $\mathcal{L}_{\textnormal{time}}=\mathcal{L}_{\textnormal{CE}}(\texttt{logits},\texttt{labels})$.
[…]
7:  Compute $\mathcal{L}_{\textnormal{total}}=(\mathcal{L}_{\textnormal{time}}+\mathcal{L}_{\textnormal{meta})}/2$.
@@
8:  Update parameters of $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$ to minimize $\mathcal{L}_{\textnormal{total}}$.
@@
Algorithm 1 summarizes one training iteration of our feature extractors $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$ as also visually depicted in Fig. 6. Given the batch of time series $X_{\textnormal{batch}}$ and metadata $C_{\textnormal{batch}}$, we randomly pick $N_{\textnormal{patch}}$ patches with horizon $L_{\textnormal{patch}}$ from each time series and metadata sample in batches $X_{\textnormal{batch}}$ and $C_{\textnormal{batch}}$ (line 1). Then, we obtain the time series and metadata embeddings for all patches through their respective feature extractors, $\phi_{\textnormal{time}}$ for time series and $\phi_{\textnormal{meta}}$ for metadata (line 2). Finally, we compute the dot product of time series and metadata embeddings (line 3), and obtain the symmetric cross-entropy loss (line 5 - 7), which is used to jointly update parameters of $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$ (line 8).
@@
In essence, we learn a joint embedding space for time series and metadata by jointly training $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$. This is achieved by adjusting the feature extractors’ parameters to maximize the cosine similarity of the time series embeddings and the metadata embeddings of $N_{\textnormal{batch}}\times N_{\textnormal{patch}}$ pairs of time series and paired metadata in the batch. In our experiments, we used the Informer encoder architecture (Zhou et al., 2021) for $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$. We choose $L_{\textnormal{patch}}$ based on the length of the smallest chunk of the time series that contains metadata-specific features. We refer the readers to App. C for further details on the choices of $N_{\textnormal{patch}}$, $L_{\textnormal{patch}}$, and the encoder architecture.
Why is J-FTSD a good metric to evaluate conditional generation models?
@@ 6 Experiments
We evaluated the performance of Time Weaver across datasets featuring a diverse mix of seasonalities, discrete and continuous metadata conditions, a wide range of horizons, and multivariate correlated channels. The list of datasets and their metadata features are provided in Table 1. All models are trained on the train split, while all metrics are reported on the test split, further detailed in App. B.
Baselines: We represent the results for the CSDI and SSSD backbones for Time Weaver as Time Weaver-CSDI and Time Weaver-SSSD, respectively. Since there are no existing approaches for metadata-conditioned time series generation with categorical, continuous, and time-variant metadata features, we modify the existing state-of-the-art GAN approaches to incorporate metadata conditions, similar to Time Weaver. The GAN baselines include CNN based approaches like WaveGAN (Donahue et al., 2019), an audio-focused GAN model, and Pulse2Pulse (Thambawita et al., 2021), a model specializing in DeepFake generation. The exact training details are provided in App. D and E. We additionally tried comparing with TimeGAN (Yoon et al., 2019), a Recurrent Neural Network (RNN) based approach, and TTS-GAN (Li et al., 2022), a Transformer-based approach. However, both of these GAN models did not converge on any of the datasets. We show their training results in Section E.2.
@@
(Thambawita et al., 2021) 20.9333 0.5874 12.8096 0.6002 16.6357 0.5409 3.1805 0.6895
Time Weaver-CSDI 2.1869 0.7419 8.8352 0.8067 0.8359 0.7189 0.5465 0.7523
Time Weaver-SSSD 10.1885 0.6826 19.0671 0.49 0.4347 0.6562 1.0505 0.7456
Table 2: DM-based approaches outperform GAN-based approaches on J-FTSD and TSTR metrics. The table shows the performance of all the models (rows) on specified datasets (columns). Time Weaver-CSDI variant significantly outperforms GANs in both metrics. Time Weaver-SSSD only underperforms for the ECG dataset, but still outperforms GANs on all other datasets. Our experimental findings also confirm that lower J-FTSD scores correspond to higher AUC (TSTR) scores when tested on the original test dataset, showcasing the utility of our proposed J-FTSD metric in evaluating the quality of the generated data distribution.
Evaluation Metrics: We evaluate our approaches and the GAN baselines using the J-FTSD metric, as detailed in Sec. 5. To validate the correctness of J-FTSD’s evaluation, we also report the area under the curve (AUC)scores of a classifier trained only using synthetic data. The classifier is trained to predict the metadata given the corresponding synthetic time series. We then test this classifier on the real unseen test dataset. High accuracy indicates that our synthetic data faithfully retained critical features of the paired metadata. For the classifier, we use a standard ResNet-1D (He et al., 2016) model. We denote this metric as TSTR in Table 2. For each dataset, the categories for which we train the classifier are: Electricity - Months (12), Air Quality - Station (12), Traffic - Weather Description (11), and ECG - Heart Conditions - (71). The exact training steps of the classifiers are outlined in App. F.
@@
Does the J-FTSD metric correctly penalize when the generated time series samples are not specific to the paired metadata? In Fig. 4, we assess the sensitivity of our J-FTSD metric against previous Frechet distance-based metrics. This assessment involves introducing controlled perturbations into the time series to test the sensitivity of the metric. These perturbations include Gaussian noise- which introduces Gaussian noise of increasing variance; time warping, involving scaling adjustments; imputation- imputing the time series with local mean and label flipping- where metadata conditions are randomly changed, decoupling them from the time series. An effective metric should demonstrate an increased sensitivity when the real and generated joint distributions of time series and metadata diverge. We compare against three Frechet distance-based metrics: 1) FTSD score, which calculates the Frechet distance using only time series embeddings (derived from the $\phi_{\textnormal{time}}$ feature extractor). 2) The Context-FID score (Jeha et al., 2021), where the $\phi_{\textnormal{time}}$ feature extractor is trained to maximize similarity for similar time series. 3) The J-FTSD (Intra-Modal) score, which is calculated the same way as J-FTSD, but the time series and metadata feature extractors are trained individually to maximize the embedding similarity for similar samples. Our J-FTSD metric is the most sensitive compared to other metrics under synthetic disturbances. The key benefit of our metric can be observed in the label-flipping experiment, where only our metric increases as we increase the label-flipping probability in the paired metadata conditions. Other metrics remain unchanged and lack sufficient sensitivity because other metrics overlook paired metadata in their distance calculations, a critical factor that J-FTSD adeptly incorporates. Additionally, the J-FTSD (Intra-Modal) score remains mostly unchanged under these perturbations, showing the advantage of jointly training time series and metadata feature extractors as we do in our metric. Experiments in Fig. 4 underscore the importance of our J-FTSD metric in assessing the quality and specificity of the generated time series data.
Does the distribution of synthetic data generated by Time Weaver match the real data distribution? Across all the datasets, Time Weaver-CSDI variant consistently outperform GAN models in terms of J-FTSD scores, as shown in Table 2. Specifically, for the J-FTSD score, we beat the best GAN model by roughly 5×5$\times$5 × on the Air Quality dataset, 1.05×1.05$\times$1.05 × on the ECG dataset, 38×38$\times$38 × on the Traffic dataset, and 5×5$\times$5 × on the Electricity dataset.
Does the synthetic data generated by Time Weaver capture metadata-specific features to train an accurate classifier? When training with the generated synthetic time series data, the classifier’s accuracy in classifying metadata hinges on the presence of distinct metadata-specific features in the time series. The high TSTR scores in Table 2 strongly suggest that the data generated by Time Weaver retain the essential characteristics necessary to train classifiers that exhibit high AUC on real unseen test data. The marked improvement in TSTR scores with Time Weaver , compared to GAN models, demonstrates both the practical value and the superior quality of the synthetic data generated by our model. The Time Weaver-SSSD model fails to learn metadata-specific features only for the ECG dataset, showing subpar results for both metrics.
@@
Figure 5: Time Weaver generated time series distribution matches with the real time series distribution. Each column represents a different dataset. The real time series is in blue, while the generated time series is in red. The first & third rows correspond to the Time Weaver model, and the second & fourth rows correspond to the best-performing GAN model. The top two rows have the time series for an unseen test metadata condition, and the bottom two rows have the comparison between the frequency distributions of the real and generated time series datasets also for the unseen test metadata condition. Both results indicate that our Time Weaver model can generate realistic time series samples that are specific to the corresponding metadata condition, beating the previous state-of-the-art GAN model. In both scenarios, the GAN models fail to match the real time series and data distribution, while our Time Weaver model has learned the correct conditional distribution for the specific metadata condition, specifically for the Air Quality and Traffic datasets.
@@
Does the lower J-FTSD correlate with higher TSTR performance? The experimental data, as outlined in Table 2, exhibit a clear correlation: lower J-FTSD scores are consistently associated with higher TSTR scores on the original, unseen test dataset. This correlation is anticipated, given that both metrics evaluate the precision of the time series relative to the corresponding metadata and the closeness of the real and synthetic joint distributions. This further underscores the effectiveness of the J-FTSD metric as a reliable indicator to assess the quality of generated data.
@@
Does the synthetic data generated by Time Weaver qualitatively match the real data? Figure 5 (top two rows) displays the quality and realism of the time series data generated by the best performing Time Weaver model. This figure contrasts generated time series samples with real ones under identical metadata conditions. The comparison demonstrates that the Time Weaver model produces time series samples highly similar to real samples, effectively mapping metadata to the corresponding time series. In contrast, GAN baseline models face challenges in generating realistic time series and accurately mapping metadata. A notable example is their performance with ECG signals (2nd column): GAN models only learn to generate a noisy version of the ECG samples while our Time Weaver model generates a pristine realistic sample. We provide additional qualitative samples in App. G.
@@
Does the synthetic data generated by Time Weaver and the real data match in terms of density and spread of time series values? In Fig. 5 (bottom two rows), we extend our analysis to compare real and generated data distributions across all datasets. This is achieved by transforming real and generated time series datasets into frequency distributions over their respective values. Take, for instance, the traffic dataset: we aggregate all time series from the dataset to form a frequency distribution over their raw values for both real and generated datasets. The Time Weaver model demonstrates a significantly more accurate representation of the real time series distribution than the best performing GAN. GAN models consistently fail to learn the complex underlying distributions of real data, particularly evident in the Air Quality and Traffic datasets.

### v2

@@ 5 Joint Frechet Time Series Distance
In J-FTSD, we compute the FD between the real and generated joint distributions of time series and paired metadata. Consider samples from a real data distribution indicated by $D_{r}=\left\{(x_{1}^{r},c_{1}),\ldots,(x_{n}^{r},c_{n})\right\}$, where $x_{i}^{r}\in\mathbb{R}^{L\times F}$ indicates the time series, and $c_{i}\in\mathbb{R}^{L\times K}$ indicates the paired metadata as explained in Sec. 3. We denote the dataset of generated time series and the corresponding metadata as $D_{g}=\left\{(x_{1}^{g},c_{1}),\ldots,(x_{n}^{g},c_{n})\right\}$, where $x_{i}^{g}=G(c_{i})\ \forall\ i\in[1,n]$, and $G$ denotes any arbitrary conditional generation model, as defined in Sec. 3. First, similar to the FID and FJD computations, we project the time series and the paired metadata into a lower-dimensional embedding space using $\phi_{\textnormal{time}}(\cdot):\mathbb{R}^{L\times F}\rightarrow\mathbb{R}^{d_{\textnormal{emb}}}$ and $\phi_{\textnormal{meta}}(\cdot):\mathbb{R}^{L\times K}\rightarrow\mathbb{R}^{d_{\textnormal{emb}}}$ as the respective feature extractors, where $d_{\textnormal{emb}}$ is the size of the embedding. We concatenate these time series and metadata embeddings to create a joint embedding space. We then calculate the FD over the joint embedding space. As such, the J-FTSD metric is formally defined as:
@@
$z^{d}_{i}$ $=\phi_{\textnormal{time}}(x^{d}_{i})\oplus\phi_{\textnormal{meta}}(c_{i})\qquad\qquad\forall i:(x_{i}^{d},c_{i})\in D_{d},$
@@
In essence, J-FTSD computes the FD between the Gaussian approximations of the real and generated joint embedding datasets. In Eq. 6, $\mu_{z^{r}}$, $\mu_{z^{g}}$ and $\Sigma_{z^{r}}$, $\Sigma_{z^{g}}$ are the mean and the variance of the Gaussian approximations of the real and generated joint embedding datasets, respectively.
@@ 12 Stations, 5 years, 12 months,
Training Feature Extractors: Now, we describe our approach to obtain the feature extractors $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$. As explained in Sec. 2, DeVries et al. (2019) suggest using separate encoders for data samples and conditions. However, they only deal with a specific type of condition, and this naturally poses a problem for a straightforward extension of their approach to our case, where the metadata could be any arbitrary combination of categorical, continuous, and time-varying features. As such, we propose a novel approach to train the feature extractors $\phi_{\textnormal{meta}}$ and $\phi_{\textnormal{time}}$ specific to the time series domain. We jointly train $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$ with contrastive learning to better capture the joint distribution of the time series and paired metadata as contrastive learning is a commonly used method to map data from various modalities into a shared latent space (Yuan et al., 2021; Zhang et al., 2022; Ramesh et al., 2022).
Figure 4: J-FTSD correctly penalizes the conditional time series data distribution. A good metric should penalize the conditional generation approaches for not being specific to the metadata and deviating from real time series data distribution. As such, we compare the sensitivity of different distance metrics under various synthetic disturbances on the Air Quality dataset (starting from the left); we add Gaussian noise, warp, impute, and randomly change the metadata of the time series samples. We clearly show that as the amount of perturbation increases, our J-FTSD metric (in red) shows the highest sensitivity, correctly capturing the dissimilarities between the perturbed and the original datasets. In contrast, the other metrics remain unchanged or show lower sensitivity.
@@
1:  Input: Time series feature extractor $\phi_{\textnormal{time}}$, Metadata feature extractor $\phi_{\textnormal{meta}}$, Time series batch $X_{\textnormal{batch}}$, Paired Metadata batch $C_{\textnormal{batch}}$, Number of patches $N_{\textnormal{patch}}$, Patch length $L_{\textnormal{patch}}$, Batch size $N_{\textnormal{batch}}$.
@@
2:  Randomly select $N_{\textnormal{patch}}$ patches of length $L_{\textnormal{patch}}$ from each sample in $X_{\textnormal{batch}}$ and $C_{\textnormal{batch}}$ to generate $X_{\textnormal{batch}}^{\textnormal{patch}}$ and $C_{\textnormal{batch}}^{\textnormal{patch}}$.
@@
3:  Obtain the time series and metadata embedding - $\phi_{\textnormal{time}}(X_{\textnormal{batch}}^{\textnormal{patch}})$ and $\phi_{\textnormal{meta}}(C_{\textnormal{batch}}^{\textnormal{patch}})$ respectively.
4:  Obtain the logits - $\phi_{\textnormal{time}}(X_{\textnormal{batch}}^{\textnormal{patch}})^{T}\phi_{\textnormal{meta}}(C_{\textnormal{batch}}^{\textnormal{patch}})$.
5:  Define the labels - $0,1,2,\dots,N_{\textnormal{batch}}\times N_{\textnormal{patch}}-1]$.
6:  Compute $\mathcal{L}_{\textnormal{time}}=\mathcal{L}_{\textnormal{CE}}(\texttt{logits},\texttt{labels})$.
7:  Compute $\mathcal{L}_{\textnormal{meta}}=\mathcal{L}_{\textnormal{CE}}(\texttt{logits.T},\texttt{labels})$.
8:  Compute $\mathcal{L}_{\textnormal{total}}=(\mathcal{L}_{\textnormal{time}}+\mathcal{L}_{\textnormal{meta})}/2$.
@@
9:  Update parameters of $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$ to minimize $\mathcal{L}_{\textnormal{total}}$.
@@
Algorithm 1 summarizes one training iteration of our feature extractors $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$. This is visually depicted in Fig. 6. Given the batch of time series $X_{\textnormal{batch}}$ and metadata $C_{\textnormal{batch}}$, we randomly pick $N_{\textnormal{patch}}$ patches with horizon $L_{\textnormal{patch}}$ from each time series and metadata sample in batches $X_{\textnormal{batch}}$ and $C_{\textnormal{batch}}$ (line 1). Then, we obtain the time series and metadata embeddings for all patches through their respective feature extractors, $\phi_{\textnormal{time}}$ for time series, and $\phi_{\textnormal{meta}}$ for metadata (line 2). Finally, we compute the dot product of time series and metadata embeddings (line 3) and obtain the symmetric cross-entropy loss (lines 4 - 7), which is used to jointly update the parameters of $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$ (line 8).
@@
In essence, we learn a joint embedding space for time series and metadata by jointly training $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$. This is achieved by adjusting the feature extractors’ parameters to maximize the cosine similarity of the time series embeddings and the metadata embeddings of $N_{\textnormal{batch}}\times N_{\textnormal{patch}}$ pairs of time series and paired metadata in the batch. In our experiments, we used the Informer encoder architecture (Zhou et al., 2021) for $\phi_{\textnormal{time}}$ and $\phi_{\textnormal{meta}}$. We choose $L_{\textnormal{patch}}$ based on the length of the smallest chunk of the time series that contains metadata-specific features. We refer the readers to Section A.3 for further details on the choices of $N_{\textnormal{patch}}$, $L_{\textnormal{patch}}$, and the encoder architecture.
Why is J-FTSD a good metric for evaluating conditional generation models?
@@ 6 Experiments
We evaluated the performance of Time Weaver across datasets featuring a diverse mix of seasonalities, discrete and continuous metadata conditions, a wide range of horizons, and multivariate correlated channels. The list of datasets and their metadata features are provided in Table 1. All models are trained on the train split, while all metrics are reported on the test split, further detailed in Section A.2.
@@
[…]
Table 2: Time Weaver outperforms GAN-based approaches on the J-FTSD and TSTR metrics. The table shows the performance of all the models (rows) on specified datasets (columns). The Time Weaver variants significantly outperform GANs on the J-FTSD and TSTR metrics. Our experimental findings also confirm that lower J-FTSD scores correspond to higher AUC (TSTR) scores when tested on the original test dataset, showcasing the utility of our proposed J-FTSD metric in evaluating the quality of the generated data distribution. We report the mean and standard deviation of both metrics, averaged over three seeds.
Baselines: We represent the results for the CSDI and SSSD backbones for Time Weaver as Time Weaver-CSDI and Time Weaver-SSSD, respectively. Since there are no existing approaches for metadata-conditioned time series generation with categorical, continuous, and time-variant metadata features, we modify the existing state-of-the-art GAN approaches to incorporate metadata conditions, similar to Time Weaver. The GAN baselines include CNN-based approaches like WaveGAN (Donahue et al., 2019), an audio-focused GAN model, and Pulse2Pulse (Thambawita et al., 2021), a model specializing in DeepFake generation. Additionally, we compare the Time Weaver variants against a diffusion model with a U-Net 1D (Ronneberger et al., 2015) denoiser backbone. We chose the 1D variant of U-Net as the basic U-Net architecture is the most preferred denoiser backbone used in pixel-space image diffusion. The exact training details are provided in Sections A.4 and A.5. We additionally tried comparing with TimeGAN (Yoon et al., 2019), a Recurrent Neural Network (RNN) based approach, and TTS-GAN (Li et al., 2022), a Transformer-based approach. However, both of these GAN models did not converge on any of the datasets. We show their training results in Section A.5.4.
Evaluation Metrics: We evaluate our approaches and the GAN baselines using the J-FTSD metric, as detailed in Sec. 5. To validate the correctness of J-FTSD’s evaluation, we also report the area under the curve (AUC) scores of a classifier trained only using synthetic data. The classifier is trained to predict the metadata given the corresponding synthetic time series. We then test this classifier on the real unseen test dataset. High accuracy indicates that our synthetic data faithfully retains critical features of the paired metadata. We use a standard ResNet-1D (He et al., 2016) model for the classifier. We denote this metric as TSTR in Table 2. For each dataset, the categories for which we train the classifier are listed as follows: Electricity - Months (12), Air Quality - Stations (12), Traffic - Weather Descriptions (11), and ECG - Heart Conditions (71). The exact training steps of the classifiers and a detailed description of the metrics are outlined in Section A.6.
@@
Does the J-FTSD metric correctly penalize when the generated time series samples are not specific to the paired metadata?
@@
Figure 5: Time Weaver generated time series distributions match the real time series distributions. Each column represents a different dataset. The real time series is in blue, while the generated time series is in red. The first & third rows correspond to the Time Weaver model, and the second & fourth rows correspond to the best-performing GAN model. The top two rows have the real and generated time series for unseen test metadata conditions. The bottom two rows compare the histograms of the real and generated time series values aggregated over their respective datasets, also for unseen test metadata conditions. Both results indicate that our Time Weaver model can generate realistic time series samples that are specific to the corresponding metadata conditions, beating the previous state-of-the-art GAN model. In both scenarios, the GAN model fails to match the real time series data distribution, while our Time Weaver model has learned the correct conditional distribution for specific metadata conditions, specifically for the Air Quality and Traffic datasets.
In Fig. 4, we assess the sensitivity of our J-FTSD metric against previous FD-based metrics. This assessment involves introducing controlled perturbations into the time series to test the sensitivity of the metric. These perturbations include Gaussian noise - which introduces Gaussian noise of increasing variance; time warping, involving scaling adjustments; imputation - imputing the time series with local mean and label flipping - where metadata conditions are randomly changed, decoupling them from the time series. An effective metric should demonstrate an increased sensitivity when the real and generated joint distributions of time series and metadata diverge. We compare J-FTSD against three FD-based metrics: 1) the FTSD score, which calculates the FD using only time series embeddings (derived from the $\phi_{\textnormal{time}}$ feature extractor); 2) the Context-FID score (Jeha et al., 2021) where $\phi_{\textnormal{time}}$ is trained to maximize similarity of similar time series; 3) the J-FTSD (Intra-Modal) score, which differs from J-FTSD in that the time series and metadata feature extractors are trained individually to maximize the embedding similarity for similar samples. Our J-FTSD metric is the most sensitive when compared to other metrics under synthetic disturbances. The key benefit of our metric can be observed in the label-flipping experiment, where only our metric increases as we increase the label-flipping probability in the paired metadata conditions. Other metrics remain unchanged and lack sufficient sensitivity because they overlook paired metadata in their distance calculations, a critical factor that J-FTSD adeptly incorporates. Additionally, the J-FTSD (Intra-Modal) score remains mostly unchanged under these perturbations, highlighting the advantage of the joint training of the time series and metadata feature extractors in our metric. Experiments in Fig. 4 underscore the importance of our J-FTSD metric in assessing the quality of the generated time series data.
How does Time Weaver compare against other approaches on real-world datasets? Across all the datasets, the Time Weaver variants consistently outperform GAN models in terms of J-FTSD scores, as shown in Table 2. Specifically, we beat the best GAN model by roughly $6\times$ on the Air Quality dataset, $1.75\times$ on the ECG dataset, $4\times$ on the Electricity dataset, and more than $40\times$ on the Traffic dataset. To understand this performance gain, we conduct an ablation study where we compare the Time Weaver-CSDI model and WaveGAN (Donahue et al., 2019) in the presence and absence of metadata on the Air Quality and Electricity datasets. We use the FTSD metric, which is the Frechet Distance without considering the metadata embeddings, as our evaluation metric. The quantitative results of this study are summarized in Table 8. We find that Time Weaver provides lower FTSD for the Electricity dataset, and WaveGAN provides lower FTSD for the Air Quality dataset in the absence of metadata. However, in the presence of metadata, we note that Time Weaver outperforms WaveGAN on both datasets. The key insight is that Time Weaver processes metadata more efficiently than GAN-based approaches to generate metadata-specific time series. This observation aligns with the recent work (Ding et al., 2020) that shows GAN s perform poorly with continuous conditions. The metadata conditions in the time series domain are so complex (a mix of categorical, continuous, and time-varying conditions) that the GAN models fail to learn a good conditional time series distribution.
Additionally, we compare Time Weaver against a diffusion model baseline with U-Net 1D (Ronneberger et al., 2015) as the denoiser backbone. The quantitative results for this comparison are provided in Table 9. We note that the best-performing Time Weaver model outperforms the diffusion model baseline with the U-Net 1D denoiser, with an average improvement of 9% in TSTR and 17% in J-FTSD metrics. This superior performance highlights the importance of time series-specific innovations in Time Weaver for complex metadata-specific time series generation.
Does the synthetic data generated by Time Weaver capture metadata-specific features to train an accurate classifier? When training with the generated synthetic time series data, the classifier’s accuracy in classifying metadata hinges on the presence of distinct metadata-specific features in the time series. The high TSTR scores in Table 2 strongly suggest that the data generated by Time Weaver retain the essential characteristics necessary to train classifiers that exhibit high AUC on real unseen test data. The marked improvement in TSTR scores with Time Weaver when compared to GAN models demonstrates both the practical value and the superior quality of the synthetic data generated by our model.
@@
Does lower J-FTSD correlate with higher TSTR performance? The experimental data, as outlined in Table 2, exhibit a clear correlation: lower J-FTSD scores are consistently associated with higher TSTR scores on the original, unseen test dataset. This correlation is expected as both metrics evaluate the precision of the generated time series relative to the corresponding metadata, as well as the closeness of the real and generated joint distributions. This further underscores the effectiveness of the J-FTSD metric as a reliable indicator to assess the quality of generated data. We further solidify the validity of J-FTSD by showing that lower values of J-FTSD correspond to lower values of the Dynamic Time Warping (DTW) metric, where DTW is computed between the real and generated time series samples. Lower DTW indicates higher similarity between the real and generated time series. We refer the reader to Section A.7.3 for further details and experimental results.
@@
Does the synthetic data generated by Time Weaver qualitatively match the real data? Figure 5 (top two rows) displays the quality and realism of the time series data generated by the best-performing Time Weaver model. This figure contrasts generated time series samples with real ones under identical metadata conditions. The comparison demonstrates that the Time Weaver model produces time series samples that are highly similar to real samples, effectively mapping metadata to the corresponding time series. In contrast, GAN baseline models face challenges in generating realistic time series and accurately mapping metadata. A notable example is their performance with ECG signals (2nd column): GAN models only learn to generate a noisy version of the ECG samples while our Time Weaver model generates a pristine, realistic sample. We provide additional qualitative samples in Section A.11.
@@
Does the synthetic data generated by Time Weaver match the real data in terms of density and spread of time series values? In Fig. 5 (bottom two rows), we extend our analysis to compare real and generated data distributions across all datasets. This is achieved by transforming the real and generated time series datasets into histograms of their respective values. Take, for instance, the traffic dataset: we aggregate all samples to form a histogram over the raw time series values for both real and generated datasets. The Time Weaver model provides a significantly more accurate representation of the real time series distribution than the best-performing GAN model. GAN models consistently fail to learn the complex underlying distributions of real data, particularly evident in the Air Quality and Traffic datasets.
Does Time Weaver retain the causal relationship between the input metadata and the corresponding generated time series? We analyze Time Weaver’s ability to retain causal relationships using the Air Quality dataset. We note that there exist physical models that indicate the effect on the particulate matter level (one of the time series channels) whenever there is rainfall (metadata input). We show through qualitative examples in Fig. 11 that the generated particulate matter levels adhere to this effect. However, we note that this is not a rigorous proof of causality. We aim to build upon this work to develop time series generative models that can maintain causal relationships between the input metadata and the generated time series.
Training and Inference Complexity: Here, we discuss the training and inference complexity of our approach. For the inference complexity, we make the following observations:
  • The Time Weaver-CSDI denoiser’s inference time complexity reflects that of any transformer architecture, scaling quadratically with the time series length $L$ and linearly with respect to the hidden dimension size $H$ and the number of diffusion steps $T$.
  • The Time Weaver-SSSD denoiser’s inference time complexity reflects that of the S4 model (Alcaraz & Strodthoff, 2022), scaling quadratically with the hidden dimension size $H$ and the number of diffusion steps $T$.
  • Our metadata encoder’s inference time complexity is the same as that of the transformer architecture, $\mathcal{O}(L^{2}K+K^{2}L)$, where $K$ is the number of metadata conditions. As the metadata encoder only runs once for a single sample generation, its time complexity is not affected by the number of diffusion steps.
Additionally, we tabulate the exact time taken to generate a sample for all datasets in Table 11. For training complexity, we show the rate of change of J-FTSD for Time Weaver and GAN models. Specifically, we considered the Air Quality dataset, and the results are shown in Fig. 10. As expected, we observe that the rate of decrease of J-FTSD is quicker for Time Weaver when compared to GAN models.